OpenStudy (anonymous):

A student wrote that a number pattern such that each number is 3 more than 4 times the previous number. What is the sixth term in the pattern if the first term is 6?

OpenStudy (jhannybean):

\[3+4(n-1) \] where \[a_1 = 6\]

OpenStudy (jack1):

the 6th term in the series using that formula would be 23... assuming n=a1=6... yeah?

OpenStudy (jack1):

is that right?

OpenStudy (jhannybean):

n represents many numbers, so if we take the first number, 6, you would just plug that in, i believe

OpenStudy (jack1):

im not sure that's right... the 2nd term should b 27 in the series, if the first term is 6

OpenStudy (jhannybean):

How would the equation be written for the sequence?

OpenStudy (hunus):

http://labs.codecademy.com/BBcN#:workspace Here's a cool little program that will show you the numbers in the sequence

OpenStudy (hunus):

\[\Large a_{n}=4a_{n-1} + 3\] Where \[a_1 = 6\]

OpenStudy (jhannybean):

Oh, you're right, Jack1. Since this is a pattern of adding +4 to every number in the sequence, we consider this an arithmetic sequence. Hunus wrote the equation out better :P

OpenStudy (jack1):

1st term 6 2nd term 27 3rd term 111 4th term 447 ... have i got the right pattern here...?

OpenStudy (hunus):

Yup, look at the link I posted

OpenStudy (jhannybean):

I'm not entirely sure what to input in there, and it's a bit confusing :P

OpenStudy (hunus):

Ah, just press the run button at the top and it will ask you which term you want it to compute up to

OpenStudy (jack1):

@Hunus can you explain your equation there? is that to the power of anything?

OpenStudy (hunus):

Nope, it says the nth term of a sequence\[\Large a_n\]is three plus the term before that one \[\Large a_{n-1}\] multiplied by four

OpenStudy (jack1):

ok, so how does your equation take into account the previous term...? sorry, maybe i'm reading it wrong, but to me that equation looks like: a(n) = [4 x a x (n-1) ] +3 ...is that right?

OpenStudy (jack1):

and a = 6

OpenStudy (hunus):

The previous term is a_(n-1)

OpenStudy (hunus):

a_1 = 6 a_2 = 4*a_1 + 3 = 4*6+3

OpenStudy (jack1):

but that only lets you compute them sequentially, as you need the previous term

OpenStudy (jack1):

it's not really an equation that lets you work out any term of a series

OpenStudy (jack1):

yeah...? or am i still not understanding?

OpenStudy (hunus):

You're correct, you would have to do them one by one

OpenStudy (jack1):

ok, sorry it wasn't a put down, i'm here for the knowledge, and i thought I was missing something

OpenStudy (jack1):

what about for the whole series: term (n) in the series = ( 1st term in the series x 4^(n-1) ) + ( (4^(n-1) -1) \[a _{n} = \left\{ a _{1} \times4^{(n-1)}\right\} + \left[ 4^{(n-1)} -1\right]\]

OpenStudy (jack1):

so the first 6 terms are 6, 27, 111, 447, 1791, 7167... and now all you have to do is plug in a1 = 6 and which term you want (n)...

OpenStudy (hunus):

Yup